OpenStudy (anonymous):

find the exxact value of the following: cos(105degrees)

OpenStudy (anonymous):

find 105 degrees on the unit circle of the attached cheat sheet the first coordinate is cosine

OpenStudy (anonymous):

ok that doesn't work, does it we have to do something else

OpenStudy (anonymous):

recognizing that \[105=60+45\] you can use the addition angle formula for cosine

OpenStudy (anonymous):

\[\cos(105)=\cos(60+45)\\ \cos(60+45)=\cos(60)+\cos(45)-\sin(60)\sin(45)\]

OpenStudy (anonymous):

where did you ge that formula